OpenStudy (anonymous):

write an equation of the line containing the specified point and perpendicular to the indicated line (0,6),2x -5=y

OpenStudy (anonymous):

Ok so all you need is a point and a slope. You already have your point (0,6) and now all you need is the slope which you can get from this line: 2x -5=y

OpenStudy (anonymous):

Can you tell me the slope of the line 2x -5=y

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

where m is the slope. Can you give it one more shot and tell me the slope?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

great but since the line we are looking for is perpendicular to this the slope actually has to be a negative reciprocal. That's fancy math talk of saying - 1/2

OpenStudy (anonymous):

now we use the point slope form. \[y-y_1=m(x-x_1)\] where \[(x_1,y_1)=(0,6)\]\[m=-\frac{1}{2}\]

OpenStudy (anonymous):

plug that in and solve for y if you have to.

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

y= -1/2 x-6

OpenStudy (anonymous):

y= -1/2 x+6 actually because you add 6 to both sides. But close enough good job!

OpenStudy (anonymous):

thank you